$SPAD/src/input bop.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool bop.output
)set message test on
)set message auto off
)clear all
--S 1 of 17
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 1
--S 2 of 17
deq := D(y x, x, 2) + D(y x, x) + y x = 0
--R
--R
--R         ,,       ,
--R   (2)  y  (x) + y (x) + y(x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 2
--S 3 of 17
nary? y
--R
--R
--R   (3)  true
--R                                                                Type: Boolean
--E 3
--S 4 of 17
unary? y
--R
--R
--R   (4)  false
--R                                                                Type: Boolean
--E 4

--S 5 of 17
opOne := operator('opOne, 1)
--R
--R
--R   (5)  opOne
--R                                                          Type: BasicOperator
--E 5
--S 6 of 17
nary? opOne
--R
--R
--R   (6)  false
--R                                                                Type: Boolean
--E 6
--S 7 of 17
unary? opOne
--R
--R
--R   (7)  true
--R                                                                Type: Boolean
--E 7
--S 8 of 17
arity opOne
--R
--R
--R   (8)  1
--R                                          Type: Union(NonNegativeInteger,...)
--E 8
--S 9 of 17
name opOne
--R
--R
--R   (9)  opOne
--R                                                                 Type: Symbol
--E 9
--S 10 of 17
is?(opOne, 'z2)
--R
--R
--R   (10)  false
--R                                                                Type: Boolean
--E 10
--S 11 of 17
is?(opOne, "opOne")

--R
--R
--R   (11)  true
--R                                                                Type: Boolean
--E 11
--S 12 of 17
properties y
--R
--R
--R   (12)  table()
--R                                           Type: AssociationList(String,None)
--E 12
--S 13 of 17
setProperty(y, "use", "unknown function" :: None )
--R
--R
--R   (13)  y
--R                                                          Type: BasicOperator
--E 13
--S 14 of 17
properties y
--R
--R
--R   (14)  table("use"= NONE)
--R                                           Type: AssociationList(String,None)
--E 14
--S 15 of 17
property(y, "use") :: None pretend String
--R
--R
--R   (15)  "unknown function"
--R                                                                 Type: String
--E 15
--S 16 of 17
deleteProperty!(y, "use")
--R
--R
--R   (16)  y
--R                                                          Type: BasicOperator
--E 16
--S 17 of 17
properties y
--R
--R

--R   (17)  table()
--R                                           Type: AssociationList(String,None)
--E 17
)spool
)lisp (bye)

———-

References^{[1] nothing}

7